a. M1= your mass in kg

b. M2= any body or any object and

determine the mass in kg.

C. R = 2.2 meters


I. Solve for Fg.

I. R is constant, Mjis doubled,

Mais constant, Find Fe, Will the

attraction become stronger or

weaker?

Ill. M is constant, Me is constant,

R is doubled, Solve for Fg . Will it

become weaker or stronger?

IV. M1 is constant, My is constant,

R is reduced to half. What is Fe?

Will it become stronger or

weaker?


1
Expert's answer
2021-12-12T16:42:48-0500

I. For 74-kg me and a 300-g calculator at a distance of 60 cm:


"F=G\frac{M_1M_2}{R^2}=4.12\text{ nN}."


II. The force will increase fourfold if both masses are doubled:


"F_M=G\frac{(2M_1)(2M_2)}{R^2}=4F."

III. The force will decrease fourfold:


"F_R=G\frac{(M_1)(M_2)}{(2R)^2}=\frac F4."


IV. The force will increase fourfold:


"F_E=G\frac{(M_1)(M_2)}{(R/2)^2}=4F."